# A Comprehensive Benchmark for Electrocardiogram Time-Series

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27-31, 2025; Dublin, lrelandsubmissionid: 332ccs: Computing methodologiesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
[…]
Inspired by the success of large language models, pre-training large time-series models (LTMs) has become a mainstream trend in time-series modeling (Liu et al., 2024; Gao et al., 2024). As a time-series data type, ECG is often integrated with other time-series datasets, such as meteorological and financial data, to pre-train LTMs (Ma et al., 2024). However, this integration overlooks the unique characteristics of ECG, leading to potential misunderstandings. Compared to other time-series data, ECG has the following features: (1) It exhibits quasi-periodic behavior, which differs from other time-series with no periodicity (e.g., financial data) or stable periodicity (e.g., weather data). This is because the periodicity of ECG is influenced by physiological factors, such as heart rate and myocardial ischemia (Wagner and Strauss, 2014; Chakraborty et al., 2018); (2) ECG is applied in a wide range of complex applications, such as disease diagnosis based on classification and maternal-fetal waveform prediction based on generations. These tasks differ from the primarily predictive goals of other time-series data, like stock price forecasting (Liu et al., 2024; Zhou et al., 2022). Therefore, it is necessary to construct a specialized benchmark for ECG to validate existing methods and guide the design of new ones.
In this paper, we propose a comprehensive benchmark to standardize ECG learning and evaluation, comprising the following three parts:
@@
    We propose an ECG benchmark comprising: (a) comprehensive multi-task ECG evaluations, integrating classification, detection, forecasting, and generation to represent the underlying applications of ECG; (b) a novel metric, Feature-based Fréchet Distance (FFD), as a supplement to MSE, to assess semantic fidelity of ECG signals; (c) a novel architecture, Patch Step-by-Step Model (PSSM), specifically tailored for ECG, which hierarchically encodes ECG signals through adaptive patching.
@@ 2. Related Works
ECG Analysis Methods. Current ECG analysis can be categorized into two main camps: (1) Traditional signal processing-based methods. For example, researchers used Fourier analysis and wavelet transform for QRS complex detection (Murray et al., 1980; Kumar et al., 2019) and applied RR interval variability for arrhythmia classification (Tsipouras et al., 2002). (2) Deep learning-based methods, where researchers applied deep models to significantly improve the performance of ECG tasks. For example, they used Tranformer-based frameworks for disease classification (El-Ghaish and Eldele, 2024; Miotto et al., 2018; Liu et al., 2021), CNN-based frameworks for waveforms detection (Peimankar and Puthusserypady, 2021), and GAN-based frameworks for ECG generation (Sarkar and Etemad, 2021; Mohebbian et al., 2021). However, these methods are designed and validated for specific ECG tasks and cannot fully capture the complex characteristics of ECG signals. In this paper, we propose a comprehensive ECG benchmark that thoroughly evaluates the ability of time-series models to address a wide variety of ECG applications.
@@
[…]
Time Series Foundational Models. Time-series analysis has evolved from classical statistical methods like autoregressive integrated moving average model (Box et al., 2015), which models temporal dependencies through autoregression and moving averages, to deep learning approaches such as recurrent neural networks (Rumelhart et al., 1986), long short-term memory model (Hochreiter, 1997), and transformers (Vaswani, 2017), which capture long-range dependencies with recurrent or self-attention mechanisms. Recent advancements have focused on improving transformer-based architectures for time-series tasks: Informer (Zhou et al., 2021) introduced linear complexity sparse attention mechanisms for long sequences, while Medformer (Wang et al., 2024) tailored transformers for medical time-series featuring via multi-scale data fusion.
Large Time-Series Models (LTMs). LTMs can be categorized into two main approaches: (1) Direct application of large language models (LLMs) to time-series modelling. For example, OneFitsAll (Zhou et al., 2023) leveraged GPT-2 for time-series forecasting, and Time-LLM (Jin et al., 2023) enhanced zero-shot generalization of LLMs through prompt engineering. (2) Pre-training models on large-scale time-series datasets (Dooley et al., 2024; Ekambaram et al., 2023, 2024). For example, Timer (Liu et al., 2024) pre-trained a Transformer decoder using next-token prediction for time-series forecasting, and UniTS (Gao et al., 2024) extended Transformer encoder with multi-task heads and jointly pre-trained on multimodal time-series data using a masked reconstruction objective. Furthermore, ECGPT (Davies et al., 2024) pre-trained a Transformer decoder on ECG data, while its small pre-training dataset and singular focus limited its ability to address the diverse downstream applications of ECG. In this paper, we highlight the unique characteristics of ECG compared to other time-series data and argue that LTMs may not necessarily be suitable for various ECG applications. Our extensive experiments demonstrate our benchmark’s rationality, our metric’s robustness, and our model’s effectiveness, offering new insights for the development of ECG analysis.
@@ 3.1. Preliminary: Time-series Models
where $s_{i+1}\!\in\!\mathbb{R}^{1\times m}$ is the ground truth for the $(i\!+\!1)$-th time-series tokens with $m$ time length, $e_{i}\!\in\!\mathbb{R}^{1\times d}$ denotes the features of tokens at $s_{i}$ extracted by the time-series model, $d$ denotes the hidden dimension, and $w\in\mathbb{R}^{d\times m}$ is a linear layer aligning the hidden features with the prediction space.
@@ 3.3. ECG Metric: Feature-based Fréchet Distance
$\tilde{\mu}=\frac{1}{N}\sum_{i=1}^{N}\bm{e}_{i},\quad\tilde{\Sigma}=\frac{1}{N-1}\sum_{i=1}^{N}(\bm{e}_{i}-\tilde{\mu})^{\top}(\bm{e}_{i}-\tilde{\mu})$,
@@ 3.4. ECG Model: Patch Step-by-Step Model
$\bm{e}^{1}&=\text{Embeding}(\bm{x}),\\ \bm{e}^{i+1}&=\text{ConvBlock}^{i}(\text{Patch}(\bm{e}^{i})),\\ \end{aligned}$
@@
$\bm{e}^{i+1}&=\text{ConvBlock}^{i}({\text{UnPatch}(\bm{e}^{i})}),\\ \bm{e}&=\bm{e}^{2l+1}w,\end{aligned}$
where $i\!=\!l\!+\!1,\dots,2l$ denotes $l$ decoder layers, $w$ is a linear layer and $\bm{e}$ denotes the final extracted features for ECG. The UnPatch layer restores $\bm{e}^{i+1}$ by weighted copying the features from the last layer and then fed into ConvBlock: $(e^{i+1}_{2k-1},e^{i+1}_{2k})$$=(\text{ConvBlock}^{i}(c_{1}e^{i}_{k}),$ $\text{ConvBlock}^{i}({c_{2}e^{i}_{k}}))$, where $c_{1}$ and $c_{2}$ are learnable parameters, and ConvBlock halves the hidden dimension.
@@ 4.1. Dataset
  • CPSC2018 contains 7 disease classes, including ST-segment depression, etc.
@@
  • Fetal ECG with Pregnancy and Labor (FEPL) (Matonia et al., 2020) collects direct fetal ECGs and maternal abdominal ECGs annotated with fetal R-waves.
@@
  • MIT-BIH Arrhythmia Database (MITDB, MITPDB) (Moody and Mark, 2001; Maršánová et al., 2018), MITDB contains 48 ECG records and annotates the QRS complex, and MITPDB complements MITDB by adding P waves annotations.
@@
  • PTB Diagnostic ECG Database (PTB, PTBXL) (Bousseljot et al., 1995; Wagner et al., 2020), PTB contains 549 records from 290 subjects. PTBXL ECG contains 21,799 clinical 12-lead ECGs from 18,869 patients, with each ECG lasting 10 seconds. They are often used in ECG classification tasks.
@@
We further organized an extension ECG Dataset for LTMs pre-training, comprising AF, CPSC2018, SPB, PTB, PTBXL, NFE, MIMICSub, The Georgia 12-lead ECG Challenge Database (Goldberger et al., 2000), and the Shaoxing and Ningbo Hospital ECG Database (Zheng et al., 2020b, a). This dataset contains 98,513 subjects processed into 8,414,834 training samples. Although the extension ECG Dataset overlaps with other datasets, the training objectives differ (e.g., AF and CPSC2018), or the overlapping samples are removed (e.g., MIMICSub). We describe more details about ECG datasets in the supplementary material.
@@ 4.2. Implementation Details
Datasets. All ECG recordings were resampled to 100Hz with 500 points fixed length. We discarded recordings with $>25\%$ missing data or abnormal morphologies, imputing missing segments via linear interpolation. Multi-channel ECG recordings were split into single-channel recordings to standardize input dimensions. The datasets were partitioned into 50% for training/fine-tuning and 50% for testing. All datasets have been stripped of sensitive information.
We addressed class imbalance by removing underrepresented classes and downsampling overrepresented ones for classification. In detection tasks with severe class imbalance (waveform annotation positions <5%), we adopted the F1-score. Non-maximum suppression was applied to eliminate redundant waveform probabilities. The predicted position must be within a $\pm$70ms window centred on the annotation position(Cai et al., 2020; Association for the Advancement of Medical Instrumentation et al., 1998). For forecasting, the predicting length is 100.
Models. We evaluated advanced LTMs introduced in Section 2: Timer (Liu et al., 2024), UniTS (Gao et al., 2024), and ECGPT (Davies et al., 2024). Both raw versions and variants are pre-trained on the extension ECG Dataset, denoted by “*”. The mapping $f$ of FFD is the LTM constructed by the transformer encoder. Additionally, we tested state-of-the-art transformer-based end-to-end models: Informer (Zhou et al., 2021) and Medformer (Wang et al., 2024) alongside our proposed PSSM.
All pre-trainings use the extension ECG Dataset and were conducted on NVIDIA RTX 4090 GPUs with a batch size of 8192 over 300 epochs. We use the Adam optimizer (Kingma and Ba, 2014) for model optimization with a learning rate of $1\times 10^{-6}$. All test experiments were conducted on NVIDIA A800 GPUs with a batch size of 1024 over 100 epochs, using the Adam optimizer with learning rates empirically selected from $\{1\!\times\!10^{-4},5\!\times\!10^{-5},1\!\times\!10^{-5}\}$. For LTMs, we freeze the backbone and only fine-tune a fully connected layer and task heads, yet remain unmodified in forecasting. Other models undergo full training. More details are provided in the supplementary material.
@@ 4.3. Analysis
Our proposed FFD is a reliable metric for assessing the quality of generated ECGs, with lower FFD values indicating better ECG fidelity. For example, as illustrated in the third row of Figure 3, PSSM accurately reconstructs the R-wave locations and amplitudes from phonocardiograms on the SST dataset, whereas other models yield clinically meaningless flat lines. This finding aligns with the quantitative results in Table 4, where PSSM achieves the best FFD as 0.216, significantly outperforming other methods, such as Timer* by 0.31 (59%). This consistency between qualitative and quantitative results is also validated in the first-row MITDB and second-row FEPL datasets.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754729isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Inspired by the success of large language models, pre-training large time-series models (LTMs) has become a mainstream trend in time-series modeling (Liu et al., 2024; Gao et al., 2024). As a time-series data type, ECG is often integrated with other time-series datasets, such as meteorological and financial data, to pre-train LTMs (Ma et al., 2024). However, this integration overlooks the unique characteristics of ECG, leading to potential misunderstandings. Compared to other time-series data, ECG has the following features: (1) It exhibits quasi-periodic behavior, which differs from other time-series with no periodicity (e.g., financial data) or stable periodicity (e.g., weather data). This is because the periodicity of ECG is influenced by physiological factors, such as myocardial ischemia (Wagner and Strauss, 2014; Chakraborty et al., 2018); (2) ECG is applied in a wide range of complex applications, such as disease diagnosis based on classification and maternal-fetal waveform prediction based on generations. These tasks differ from the primarily predictive goals of other time-series data, like stock price forecasting (Liu et al., 2024; Zhou et al., 2022). Therefore, it is necessary to construct a specialized benchmark for ECG to validate existing methods and guide the design of new ones. In this paper, we propose a comprehensive benchmark to standardize ECG research, comprising the following three parts:
@@
    We propose an ECG benchmark comprising: (a) comprehensive multi-task ECG evaluations, integrating classification, detection, forecasting, and generation to represent the underlying applications of ECG; (b) a novel metric, Feature-based Fréchet Distance (FFD), as a supplement to MSE, to assess semantic fidelity of ECG; (c) a novel architecture, Patch Step-by-Step Model (PSSM), specifically tailored for ECG, which hierarchically encodes ECG through adaptive patching.
@@ 2. Related Works
ECG Analysis Methods. Current ECG analysis can be categorized into two main camps: (1) Traditional signal processing-based methods. For example, researchers used Fourier analysis and wavelet transform for QRS complex detection (Murray et al., 1980; Kumar et al., 2019) and applied RR interval variability for arrhythmia classification (Tsipouras et al., 2002). (2) Deep learning-based methods. For example, they used Tranformer-based frameworks for disease classification (El-Ghaish and Eldele, 2024; Miotto et al., 2018; Liu et al., 2021), CNN-based frameworks for waveforms detection (Peimankar and Puthusserypady, 2021), and GAN-based frameworks for ECG generation (Sarkar and Etemad, 2021; Mohebbian et al., 2021). However, these methods are designed and validated for specific ECG tasks and cannot fully capture the complex characteristics of ECG signals. In this paper, we propose a comprehensive ECG benchmark that thoroughly evaluates the ability of time-series models to address a wide variety of ECG applications.
Time Series Foundational Models. Time-series analysis has evolved from classical statistical methods like autoregressive integrated moving average model (Box et al., 2015), to deep learning approaches such as recurrent neural networks (Rumelhart et al., 1986), long short-term memory model (Hochreiter, 1997), and transformers (Vaswani, 2017), which capture long-range dependencies with recurrent or self-attention mechanisms. Recent advancements have focused on improving transformer-based architectures for time-series tasks: Informer (Zhou et al., 2021) introduced linear complexity sparse attention mechanisms for long sequences, while Medformer (Wang et al., 2024) tailored transformers for medical time-series featuring via multi-scale data fusion.
Large Time-Series Models (LTMs). LTMs can be categorized into two main approaches: (1) Direct application of large language models (LLMs) to time-series modelling. For example, OneFitsAll (Zhou et al., 2023) leveraged GPT-2 for time-series forecasting, and Time-LLM (Jin et al., 2023) enhanced zero-shot generalization of LLMs through prompt engineering. (2) Pre-training models on large-scale time-series datasets (Dooley et al., 2024; Ekambaram et al., 2023, 2024). For example, Timer (Liu et al., 2024) pre-trained a Transformer decoder using next-token prediction for time-series forecasting, and UniTS (Gao et al., 2024) extended Transformer encoder with multi-task heads. Furthermore, ECGPT (Davies et al., 2024) pre-trained a Transformer decoder on ECG data, while its small pre-training dataset and singular focus limited its ability to address the diverse downstream applications of ECG. In this paper, we highlight the unique characteristics of ECG compared to other time-series data and argue that LTMs may not necessarily be suitable for various ECG applications. Our extensive experiments demonstrate our benchmark’s rationality, our metric’s robustness, and our model’s effectiveness, offering new insights for the development of ECG analysis.
@@ 3.1. Preliminary: Time-series Models
where $s_{i+1}\!\in\!\mathbb{R}^{1\times m}$ is the ground truth for the (i+1)(i\!+\!1)-th time-series tokens with $m$ time length, $e_{i}\!\in\!\mathbb{R}^{1\times d}$ denotes the features of tokens at $s_{i}$ extracted by the time-series model, $d$ denotes the hidden dimension, and $w\in\mathbb{R}^{d\times m}$ is a linear layer aligning the hidden features with the prediction space.
@@ 3.3. ECG Metric: Feature-based Fréchet Distance
[…]
@@ 3.4. ECG Model: Patch Step-by-Step Model
[…]
@@
$\begin{aligned} \bm{e}^{i+1}&=\text{ConvBlock}^{i}({\text{UnPatch}(\bm{e}^{i})}),\\ \bm{e}&=\bm{e}^{2l+1}w,\end{aligned}$
where $i\!=\!l\!+\!1,\dots,2l$ denotes $l$ decoder layers, $w$ is a linear layer and $\bm{e}$ denotes the final extracted features for ECG. The UnPatch layer restores $\bm{e}^{i+1}$ by weighted copying the features from the last layer and then fed into ConvBlock: $(e^{i+1}_{2k-1},e^{i+1}_{2k})$$(\text{ConvBlock}^{i}(c_{1}e^{i}_{k}),$ $\text{ConvBlock}^{i}({c_{2}e^{i}_{k}}))$, where $c_{1}$ and $c_{2}$ are learnable parameters, and ConvBlock halves the hidden dimension.
@@ 4.1. Dataset
  • CPSC2018 contains 7 disease classes.
@@
  • Fetal ECG with Pregnancy and Labor (FEPL) (Matonia et al., 2020) collects fetal ECGs and maternal ECGs annotated with fetal R-waves.
@@
  • MIT-BIH Arrhythmia Database (MITDB, MITPDB) (Moody and Mark, 2001; Maršánová et al., 2018), MITDB contains 48 ECG records, and MITPDB complements MITDB by adding P waves annotations.
@@
  • PTB Diagnostic ECG Database (PTB, PTBXL) (Bousseljot et al., 1995; Wagner et al., 2020), PTB contains 549 records from 290 subjects. PTBXL ECG contains 21,799 clinical 12-lead ECGs from 18,869 patients. They are often used in ECG classification tasks.
@@
  • European ST-T Database (EBD) (Taddei et al., 1992) contains 90 annotated Holter recordings from 79 subjects.
@@
  • Abdominal and Direct Fetal ECG (ADFECGDB) (Jezewski et al., 2012) contains maternal abdominal ECG and direct fetal ECG.
@@
We further organized an extension ECG Dataset for LTMs pre-training, comprising AF, CPSC2018, SPB, PTB, PTBXL, NFE, MIMICSub, The Georgia 12-lead ECG Challenge Database (Goldberger et al., 2000), and the Shaoxing and Ningbo Hospital ECG Database (Zheng et al., 2020; Zheng and others, 2020). This dataset contains 98,513 subjects processed into 8,414,834 training samples. Although the extension ECG Dataset overlaps with other datasets, the training objectives differ (e.g., AF and CPSC2018), or the overlapping samples are removed (e.g., MIMICSub). We describe more details about ECG datasets in the supplementary material.
@@ 4.2. Implementation Details
Datasets. All ECG recordings were resampled to 100Hz with 500 points fixed length. We discarded recordings with $25\%$ missing data or abnormal morphologies, imputing missing segments via linear interpolation. Multi-channel ECG recordings were split into single-channel recordings to standardize input dimensions. The datasets were partitioned into 50% for training/fine-tuning and 50% for testing. All datasets have been stripped of sensitive information.
We addressed class imbalance by removing underrepresented classes and downsampling overrepresented ones for classification. In detection tasks with severe class imbalance (waveform annotation positions <5%), we adopted the F1-score. Non-maximum suppression was applied to eliminate redundant waveform probabilities. The predicted position must be within a $\pm$70ms window centred on the annotation position(Cai et al., 2020; Association for the Advancement of Medical Instrumentation and others, 1998). For forecasting, the predicting length is 100.
Models. We evaluated advanced LTMs introduced in Section 2: Timer (Liu et al., 2024), UniTS (Gao et al., 2024), and ECGPT (Davies et al., 2024). Both raw versions and variants are pre-trained on the extension ECG Dataset, denoted by “*”. The mapping $f$ of FFD is the LTM constructed by the transformer encoder. Additionally, we tested state-of-the-art transformer-based end-to-end models: Informer (Zhou et al., 2021) and Medformer (Wang et al., 2024) alongside our proposed PSSM. All pre-trainings use the extension ECG Dataset and were conducted on NVIDIA RTX 4090 GPUs with a batch size of 8192 over 300 epochs. We use the Adam optimizer (Kingma and Ba, 2014) for optimization with a learning rate of $1\times 10^{-6}$. All test experiments were conducted on NVIDIA A800 GPUs with a batch size of 1024 over 100 epochs, using the Adam optimizer with learning rates empirically selected from $\{1\!\times\!10^{-4},5\!\times\!10^{-5},1\!\times\!10^{-5}\}$. For LTMs, we freeze the backbone and only fine-tune a fully connected layer and task heads, yet remain unmodified in forecasting. Other models undergo full training. More details are provided in the supplementary material.
@@ 4.3. Analysis
Our proposed FFD is a reliable metric for assessing the quality of generated ECGs, with lower FFD values indicating better ECG fidelity. As illustrated in the third row of Figure 3, PSSM accurately reconstructs the R-wave location and amplitudes from phonocardiograms on the SST dataset, whereas other models yield meaningless flat lines. This finding aligns with the quantitative results in Table 4, where PSSM achieves the best FFD as 0.216, significantly outperforming other methods, such as Timer by 0.31 (59%). This consistency between qualitative and quantitative results is also validated in the first-row MITDB and second-row FEPL datasets.